OpenStudy (anonymous):

A restaurant manager bought 20 packages of bagels. Some packages contained 6 bagels each and the rest contained 12 bagels each. There were 168 bagels in all. How many packages of 12 bagels did the manager buy?

OpenStudy (dan815):

nvm that was sparkly16 lol

OpenStudy (anonymous):

20 = (#of6packs) + (#0f12packs) 168= 6(#of6packs) + 12(#of12packs) solve those equations by substitution or elimination.

OpenStudy (anonymous):

how would I set up for example, # of 6 packs into like equation form? (:

OpenStudy (anonymous):

please help:)

OpenStudy (anonymous):

im confused....wait what r the variables representing?

OpenStudy (anonymous):

so how many packages of 12 bagels did the manager buy?

OpenStudy (anonymous):

112 is not one of the answer choices

OpenStudy (anonymous):

someone please help?

OpenStudy (anonymous):

I know the answer but not sure how to set it up as an equation 168/12 = 14 packages of 12 and 168/6 = 28 packages of 6

OpenStudy (anonymous):

i think its 8 packages of 12 bagels... because 8 (12 bagels) + 12 (6 bagels) =196 bagels in all

OpenStudy (anonymous):

14 packages of 12 bagels is not an answer choice... these are the choices by the way.. A. 6 B. 8 C. 9 D. 12

OpenStudy (anonymous):

???

OpenStudy (anonymous):

Trying to figure it out :)

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

It is 8 packages but can't figure out how to write it as an equation lol

OpenStudy (anonymous):

oh ok thats what I got to...

OpenStudy (anonymous):

i just plugged the answers back in

OpenStudy (anonymous):

:)